### v3

# On the connections between the low dimensional homology groups of $\mathit{{\rm SL}}_{2}$ and $\mathit{{\rm PSL}}_{2}$

Abstract.
In this article we study the low dimensional homology groups of the special linear group $\mathit{{\rm SL}}_{2}(A)$ and the projective special linear group $\mathit{{\rm PSL}}_{2}(A)$, $A$ a domain, through the natural surjective map $\mathit{{\rm SL}}_{2}(A)\rightarrow\mathit{{\rm PSL}}_{2}(A)$. In particular, we study the connection of the first, the second and the third homology groups of these groups over euclidean domains $\mathbb{Z}[\frac{1}{m}]$, $m$ a square free integer, and local domains.
Let $A$ be a commutative ring with unit and $\mu_{2}(A)$ be the group of $2$-roots of unity in $A$. By studying the Lyndon/Hochschild-Serre spectral sequence associated to the central extension
$1\rightarrow\mu_{2}(A)\rightarrow\mathit{{\rm SL}}_{2}(A)\rightarrow\mathit{{\rm PSL}}_{2}(A)\rightarrow 1,$
and using the homology of cyclic groups [6, pages 58-59] and the universal coefficient theorem [6, Exercise 3, §1, Chap III] it follows that, for any non-negative integer $n$, the natural homomorphism of homology groups
$H_{n}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\rightarrow H_{n}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})$
has 2-power torsion kernel and cokernel. In this article we study this map for $n\leq 3$.
As our first main theorem we prove that if $A$ is a domain, then we have the 9-term exact sequence
$0\rightarrow\mu_{2}(A)^{\sim}\rightarrow\widetilde{H}_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})\rightarrow\mu_{2}(A)\otimes_{\mathbb{Z}}H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$
$\rightarrow H_{2}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\rightarrow H_{2}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})\rightarrow\mu_{2}(A)\rightarrow H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$
$\rightarrow H_{1}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})\rightarrow 0,$
where $\widetilde{H}_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$ is a quotient of $H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$ over a certain 2-torsion subgroup and
$\mu_{2}(A)^{\sim}\simeq\begin{cases}0&\text{ if ${\rm char}(A)=2$}\\ \mathbb{Z}/4&\text{ if ${\rm char}(A)\neq 2$}\end{cases}$
(see Theorem 3.1 for the precise statement). Moreover, if ${\rm char}(A)\neq 2$, we show that the image of the map $\mu_{2}(A)^{\sim}\rightarrow\widetilde{H}_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$ coincides with the image of the map $H_{3}(\langle w\rangle,\mathbb{Z})\rightarrow\widetilde{H}_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$, where $w={\left(\!\!\begin{array}[]{cc}\!\!0&\!\!1\\ \!\!-1&\!\!0\end{array}\!\!\right)}\in\mathit{{\rm SL}}_{2}(A)$. Observe that the five term exact sequence from the right of the above exact sequence is the well-known five term exact sequence coming from the above central extension [6, Corollary 6.4, Chap. VII].
As first particular case, we study the above 9-term exact sequence over the euclidean domain $A_{m}:=\mathbb{Z}[\frac{1}{m}]$, where $m$ is a square free integer (Proposition 4.1). In particular we show that if $m$ is even, then we have the exact sequence
$\begin{array}[]{c}0\rightarrow\mathbb{Z}/4\rightarrow{\displaystyle\frac{H_{3}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})}{\mu_{2}(A_{m})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})}}\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(A_{m}),\mathbb{Z})\rightarrow 0\end{array}$
and if $m$ is odd, then the sequence
$\begin{array}[]{c}0\rightarrow\mathbb{Z}/4\rightarrow{\displaystyle\frac{H_{3}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})}{\mu_{2}(A_{m})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})}}\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(A_{m}),\mathbb{Z})\rightarrow\mathbb{Z}/2\rightarrow 0\end{array}$
is exact.
As second particular case we study the above 9-term exact sequence over local domains (Proposition 5.1). In particular, if the residue field of a local domain $A$ is different than $\mathbb{F}_{2}$, then we obtain the exact sequence
$0\rightarrow\mathbb{Z}/4\rightarrow\frac{H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}{\mu_{2}(A)\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})\rightarrow 0$
We will see that this is not necessarily true if the residue field of the local domain is $\mathbb{F}_{2}$. For example we show that for $\mathbb{Z}_{(2)}=\{a/b\in\mathbb{Q}:a,b\in\mathbb{Z},2\nmid b\}$, we have the exact sequence
$\begin{array}[]{c}0\rightarrow\mathbb{Z}/4\rightarrow{\displaystyle\frac{H_{3}(\mathit{{\rm SL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})}{\mu_{2}(\mathbb{Z}_{(2)})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})}}\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})\rightarrow\mathbb{Z}/2\rightarrow 0.\end{array}$
In all the above exact sequences the map
$\rho_{\ast}:\mu_{2}(A)\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$
is induced by the product map
$\rho:\mu_{2}(A)\times\mathit{{\rm SL}}_{2}(A)\rightarrow\mathit{{\rm SL}}_{2}(A),\ \ \ \ (a,X)\mapsto aX.$
We will show that for $A_{2}=\mathbb{Z}[\frac{1}{2}]$, $\rho_{\ast}$ is trivial and for $A_{3}=\mathbb{Z}[\frac{1}{3}]$ the map $\rho_{\ast}$ is non-trivial (see Example 4.4).
If $A$ is a local domain such that its residue field has more than two elements, then the natural product map
$\rho_{\ast}:\mu_{2}(A)\otimes_{\mathbb{Z}}H_{i}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\rightarrow H_{i+1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$
is trivial for $i=0,1$. The following question ask if this is true for $i=2$.
Question 0.1.
Let $A$ be a local domain such that $|A/\mathfrak{m}_{A}|\neq 2$. Is the product map
$\rho_{\ast}:\mu_{2}(A)\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$
trivial?
We do not know the answer of the above question even over a general infinite field. But we show that $\rho_{\ast}$ is trivial for finite fields, quadratically closed fields and real closed fields (Proposition 5.4).
Here we outline the organization of the present paper. In Section 1, we recall some results from the literature over the low dimensional homology groups of central extensions. In Section 2 we introduce the scissors congruence group, the Bloch group and formulate the Bloch-Wigner exact sequence over fields. In Section 3 we prove our 9-term exact sequence (Theorem 3.1). In Section 4, we study the above $9$-term exact sequence over the euclidean domain $\mathbb{Z}[\frac{1}{m}]$. Finally in Section 5 we investigate the above $9$-term exact sequence over local rings.
Notations. In this paper all rings are commutative, except possibly group rings, and have the unit element $1$. For a commutative ring $A$, $\mu(A)$ denotes the group of roots of unity in $A$, i.e.
$\mu(A):=\{a\in A:\text{there is $n\in\mathbb{N}$ such that $a^{n}=1$}\},$
and $\mu_{2}(A):=\{a\in A:a^{2}=1\}$. If $\mathcal{B}\rightarrow\mathcal{A}$ is a homomorphism of abelian groups, by $\mathcal{A}/\mathcal{B}$ we mean ${\rm coker}(\mathcal{B}\rightarrow\mathcal{A})$.
1. The low dimensional homology of central extensions
Let $\mathcal{A}$ be a central subgroup of a group $\mathcal{G}$. The product map $\rho:\mathcal{A}\times\mathcal{G}\rightarrow\mathcal{G}$, given by $(a,g)\mapsto ag$, induces the natural map $\rho_{\ast}:H_{3}(\mathcal{A}\times\mathcal{G},\mathbb{Z})\rightarrow H_{3}(\mathcal{G},\mathbb{Z})$. By the Künneth formula we have the exact sequence
$\begin{array}[]{c}0\rightarrow\bigoplus_{i=0}^{3}H_{i}(\mathcal{A},\mathbb{Z})\otimes_{\mathbb{Z}}H_{3-i}(\mathcal{G},\mathbb{Z})\rightarrow H_{3}(\mathcal{A}\times\mathcal{G},\mathbb{Z})\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\mathcal{A},H_{1}(\mathcal{G},\mathbb{Z}))\rightarrow 0.\end{array}$
If
$\widehat{H}_{3}(\mathcal{A}\times\mathcal{G},\mathbb{Z}):=\frac{H_{3}(\mathcal{A}\times\mathcal{G},\mathbb{Z})}{\mathcal{A}\otimes_{\mathbb{Z}}H_{2}(\mathcal{G},\mathbb{Z})\ \oplus H_{2}(\mathcal{A},\mathbb{Z})\otimes_{\mathbb{Z}}H_{1}(\mathcal{G},\mathbb{Z})},$
then from the above exact sequence we obtain the exact sequence
$0\rightarrow H_{3}(\mathcal{A},\mathbb{Z})\oplus H_{3}(\mathcal{G},\mathbb{Z})\rightarrow\widehat{H}_{3}(\mathcal{A}\times\mathcal{G},\mathbb{Z})\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\mathcal{A},H_{1}(\mathcal{G},\mathbb{Z}))\rightarrow 0$
This exact sequence splits naturally, thus we have a natural splitting
${{\rm Tor}_{1}^{\mathbb{Z}}}(\mathcal{A},H_{1}(\mathcal{G},\mathbb{Z}))\hookrightarrow\widehat{H}_{3}(\mathcal{A}\times\mathcal{G},\mathbb{Z}).$
It is straightforward to check that the image of the composite
$H_{2}(\mathcal{A},\mathbb{Z})\otimes_{\mathbb{Z}}H_{1}(\mathcal{G},\mathbb{Z})\rightarrow H_{3}(\mathcal{A}\times\mathcal{G},\mathbb{Z})\overset{\rho_{\ast}}{\longrightarrow}H_{3}(\mathcal{G},\mathbb{Z})$
is in the image of the composite
$\mathcal{A}\otimes_{\mathbb{Z}}H_{2}(\mathcal{G},\mathbb{Z})\rightarrow H_{3}(\mathcal{A}\times\mathcal{G},\mathbb{Z})\overset{\rho_{\ast}}{\longrightarrow}H_{3}(\mathcal{G},\mathbb{Z})$
(see [26, Proposition 4.4(i), Chap. V]). Thus we get the natural composite map
${{\rm Tor}_{1}^{\mathbb{Z}}}(\mathcal{A},H_{1}(\mathcal{G},\mathbb{Z}))\rightarrow\widehat{H}_{3}(\mathcal{A}\times\mathcal{G},\mathbb{Z})\rightarrow H_{3}(\mathcal{G},\mathbb{Z})/\rho_{\ast}(\mathcal{A}\otimes_{\mathbb{Z}}H_{2}(\mathcal{G},\mathbb{Z})).$
We set
$\widetilde{H}_{3}(\mathcal{G},\mathbb{Z}):=H_{3}(\mathcal{G},\mathbb{Z})/\rho_{\ast}\bigg{(}\mathcal{A}\otimes_{\mathbb{Z}}H_{2}(\mathcal{G},\mathbb{Z})\oplus{{\rm Tor}_{1}^{\mathbb{Z}}}(\mathcal{A},H_{1}(\mathcal{G},\mathbb{Z})\bigg{)}.$
In this article, for calculations purposes, we will use the (left) bar resolution $B_{\bullet}(\mathcal{G})\rightarrow\mathbb{Z}$ of $\mathbb{Z}$ over $\mathbb{Z}[\mathcal{G}]$ to study the integral homology groups $H_{n}(\mathcal{G},\mathbb{Z})$ (see [6, Chap.I, §5]):
$H_{n}(\mathcal{G},\mathbb{Z}):=H_{n}(B_{\bullet}(\mathcal{G})\otimes_{\mathbb{Z}[\mathcal{G}]}\mathbb{Z})=H_{n}(B_{\bullet}(\mathcal{G})_{\mathcal{G}}).$
If any pair of the elements $g_{1},\dots,g_{n}\in G$ commute, then we define
${\boldsymbol{c}}(g_{1},\dots,g_{n}):=\overline{\sum_{\sigma\in S_{n}}{\rm sign}(\sigma)[g_{\sigma(1)}|\dots|g_{\sigma(n)}]}\in H_{n}(\mathcal{G},\mathbb{Z}),$
where $S_{n}$ is the symmetric group of degree $n$. If $\mathcal{A}:=\langle g_{1},\dots,g_{n}\rangle\subseteq\mathcal{G}$, then the Pontryagin product (see [6, Chap. V, §5]) induces a natural map
$\begin{array}[]{c}\bigwedge_{\mathbb{Z}}^{n}\mathcal{A}\rightarrow H_{n}(\mathcal{A},\mathbb{Z})\rightarrow H_{n}(\mathcal{G},\mathbb{Z}).\end{array}$
It is straightforward to show that the above composite, maps $g_{1}\wedge\cdots\wedge g_{n}$ to ${\boldsymbol{c}}(g_{1},\dots,g_{n})$.
A function $\psi:\mathcal{A}\rightarrow\mathcal{B}$ of (additive) abelian groups is called a quadratic map if
(1) for any $a\in\mathcal{A}$, $\psi(a)=\psi(-a)$,
(2) the function $\mathcal{A}\times\mathcal{A}\rightarrow\mathcal{B}$ with $(a,b)\mapsto\psi(a+b)-\psi(a)-\psi(b)$ is bilinear.
For any abelian group $\mathcal{A}$, there is a universal quadratic map
$\gamma:\mathcal{A}\rightarrow\Gamma(\mathcal{A})$
such that for any quadratic map $\psi:\mathcal{A}\rightarrow\mathcal{B}$, there is a unique group homomorphism $\Psi:\Gamma(\mathcal{A})\rightarrow\mathcal{B}$ such that $\Psi\circ\gamma=\psi$. It is easy to see that $\Gamma$ is a functor from the category of abelian groups to itself. This functor is called Whitehead’s quadratic functor. In fact if $\mathcal{E}$ is the free abelian group generated by the symbols $w(a)$, $a\in\mathcal{A}$, then we can define $\Gamma(\mathcal{A}):=\mathcal{E}/\mathcal{K}$, where $\mathcal{K}$ is the subgroup generated by the relations
(a) $w(a)-w(-a)$=0,
(b) $w(a+b+c)-w(a+b)-w(a+c)-w(b+c)+w(a)+w(b)+w(c)=0$.
The universal quadratic map $\gamma:\mathcal{A}\rightarrow\Gamma(\mathcal{A})$ is given by $a\mapsto\gamma(a):=w(a)+\mathcal{K}$. For more on Whitehead’s quadratic functor we refer the reader to [4, §1.2].
Now let $\mathcal{A}$ be a central subgroup of $\mathcal{G}$. The inclusion $i:\mathcal{A}\rightarrow\mathcal{G}$ induces the natural map $i_{\ast}:\mathcal{A}\rightarrow H_{1}(\mathcal{G},\mathbb{Z})$, $a\mapsto{\boldsymbol{c}}(a)$. By the universal property of $\Gamma$, the quadratic map
$\psi:\mathcal{A}\rightarrow\mathcal{A}\otimes_{\mathbb{Z}}H_{1}(\mathcal{G},\mathbb{Z}),\ \ \ a\mapsto a\otimes{\boldsymbol{c}}(a),$
can be extended to a group homomorphism
(1.1) $\tau:\Gamma(\mathcal{A})\rightarrow\mathcal{A}\otimes_{\mathbb{Z}}H_{1}(\mathcal{G},\mathbb{Z}).$
[…]
is trivial too. This induces the natural map
${\rm coker}(\tau)\rightarrow H_{2}(\mathcal{G},\mathbb{Z}).$
Let $\mathcal{G}$ be a group. Let $K(\mathcal{G},n)$ be an Eilenberg-MacLane space of type $(\mathcal{G},n)$. This is a CW-complex such that
$\pi_{k}(K(\mathcal{G},n),\ast)\simeq\begin{cases}1&\text{if $k\neq n$}\\ \mathcal{G}&\text{if $k=n$}.\end{cases}$
This space is unique up to homotopy. Note that for $n\geq 2$, $\mathcal{G}$ should be abelian. It is well-known [14, Theorem 21.1] that
$\Gamma(\mathcal{A})\simeq H_{4}(K(\mathcal{A},2),\mathbb{Z}).$
Let $1\rightarrow\mathcal{A}\rightarrow\mathcal{G}\rightarrow\mathcal{Q}\rightarrow 1$ be a central extension. The standard classifying space theory gives a (homotopy theoretic) fibration [9, Chap. 6] of Eilenberg-MacLane spaces
$K(\mathcal{A},1)\rightarrow K(\mathcal{G},1)\rightarrow K(\mathcal{Q},1)$
(see [23, 5.1.28]). From this we obtain the fibration
$K(\mathcal{G},1)\rightarrow K(\mathcal{Q},1)\rightarrow K(\mathcal{A},2)$
(see [16, Lemma 3.4.2]). From the Serre spectral sequence of this fibration we obtain the sequence
[…]
$H_{4}(\mathcal{Q},\mathbb{Z})\rightarrow\ker(\tau).$
The following theorem is proved by Eckmann and Hilton in [12], which generalises Ganea’s 6-term exact sequence [26, Theorem 2.2, Chap. V].
Theorem 1.1 (Eckmann-Hilton).
For any central extension $1\rightarrow\mathcal{A}\rightarrow\mathcal{G}\rightarrow\mathcal{Q}\rightarrow 1$, we have the natural $10$-term exact sequence
$H_{4}(\mathcal{Q},\mathbb{Z})\rightarrow\ker(\tau)\rightarrow\widetilde{H}_{3}(\mathcal{G},\mathbb{Z})\rightarrow H_{3}(\mathcal{Q},\mathbb{Z})\rightarrow{\rm coker}(\tau)\rightarrow H_{2}(\mathcal{G},\mathbb{Z})\rightarrow H_{2}(\mathcal{Q},\mathbb{Z})$
$\rightarrow\mathcal{A}\rightarrow H_{1}(\mathcal{G},\mathbb{Z})\rightarrow H_{1}(\mathcal{Q},\mathbb{Z})\rightarrow 0.$
Proof.
See [12, Theorem 1.1]. ∎
Let $\mathcal{A}$ be a finite cyclic group. If $2\mid|\mathcal{A}|$, let $\mathcal{A}^{\sim}$ denote the unique non-trivial extension of $\mathcal{A}$ by $\mathbb{Z}/2$. If $2\nmid|\mathcal{A}|$, we define $\mathcal{A}^{\sim}:=\mathcal{A}$. Thus if $n=|\mathcal{A}|$, then
$\mathcal{A}^{\sim}\simeq\begin{cases}\mathbb{Z}/2n&\text{if $2\mid n$}\\ \mathbb{Z}/n&\text{if $2\nmid n$.}\end{cases}$
Let $\mathcal{A}$ be an ind-cyclic group, i.e. $\mathcal{A}$ is direct limit of its finite cyclic subgroups. Then by passing to the limit we can define $\mathcal{A}^{\sim}$. Note that if $\mathcal{A}$ has no $2$-torsion, then $\mathcal{A}^{\sim}=\mathcal{A}$. There is always a canonical injective homomorphism $\mathcal{A}\hookrightarrow\mathcal{A}^{\sim}$ whose composition with the projection $\mathcal{A}^{\sim}\rightarrow\mathcal{A}$ coincides with multiplication by $2$.
Lemma 1.2.
If $\mathcal{A}$ is an ind-cyclic group, then $\Gamma(\mathcal{A})\simeq\mathcal{A}^{\sim}$.
Proof.
This follows from the facts that the functor $\Gamma$ commutes with direct limit and
$\Gamma(\mathbb{Z}/n)\simeq\mathbb{Z}/{\rm gcd}(n^{2},2n)\simeq(\mathbb{Z}/n)^{\sim},$
where ${\rm gcd}(n^{2},2n)$ is the greatest common divisor of $n^{2}$ and $2n$ (see [4, Corollary 1.2.9]). ∎
Let $\mathcal{A}$ be an abelian group. Let $\sigma_{1}:{{\rm Tor}_{1}^{\mathbb{Z}}}(\mathcal{A},\mathcal{A})\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\mathcal{A},\mathcal{A})$ be obtained by interchanging the copies of $\mathcal{A}$. This map is induced by the involution $\mathcal{A}\otimes_{\mathbb{Z}}\mathcal{A}\rightarrow\mathcal{A}\otimes_{\mathbb{Z}}\mathcal{A}$, $a\otimes b\mapsto-b\otimes a$ [18, §2]. Let $\Sigma_{2}^{\prime}=\{1,\sigma^{\prime}\}$ be the symmetric group of order 2 and consider the following action of $\Sigma_{2}^{\prime}$ on ${{\rm Tor}_{1}^{\mathbb{Z}}}(\mathcal{A},\mathcal{A})$:
$(\sigma^{\prime},x)\mapsto-\sigma_{1}(x).$
Proposition 1.3.
Let $\mathcal{A}$ be an abelian group and $T_{\mathcal{A}}$ be its torsion subgroup. Then we have the exact sequence
$\begin{array}[]{c}0\rightarrow\bigwedge_{\mathbb{Z}}^{3}\mathcal{A}\rightarrow H_{3}(\mathcal{A},\mathbb{Z})\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(T_{\mathcal{A}},T_{\mathcal{A}})^{\Sigma_{2}^{\prime}}\rightarrow 0.\end{array}$
If $T_{\mathcal{A}}$ is an ind-cyclic group, then $\Sigma_{2}^{\prime}$ acts trivially on ${{\rm Tor}_{1}^{\mathbb{Z}}}(T_{\mathcal{A}},T_{\mathcal{A}})$ and the exact sequence
$\begin{array}[]{c}0\rightarrow\bigwedge_{\mathbb{Z}}^{3}\mathcal{A}\rightarrow H_{3}(\mathcal{A},\mathbb{Z})\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(T_{\mathcal{A}},T_{\mathcal{A}})\rightarrow 0\end{array}$
splits naturally.
Proof.
By [27, Lemma 5.5] or [5, Section §6] we have the exact sequence
$\begin{array}[]{c}0\rightarrow\bigwedge_{\mathbb{Z}}^{3}\mathcal{A}\rightarrow H_{3}(\mathcal{A},\mathbb{Z})\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\mathcal{A},\mathcal{A})^{\Sigma_{2}^{\prime}}\rightarrow 0.\end{array}$
Since ${{\rm Tor}_{1}^{\mathbb{Z}}}(\mathcal{A},\mathcal{A})\simeq{{\rm Tor}_{1}^{\mathbb{Z}}}(T_{\mathcal{A}},T_{\mathcal{A}})$, we obtain the first exact sequence. Now let $T_{\mathcal{A}}$ be an ind-cyclic group. Since ${{\rm Tor}_{1}^{\mathbb{Z}}}(\mathbb{Z}/n,\mathbb{Z}/n)\simeq\mathbb{Z}/n$, the action of $\Sigma_{2}^{\prime}$ on ${{\rm Tor}_{1}^{\mathbb{Z}}}(\mathbb{Z}/n,\mathbb{Z}/n)$ is trivial. Now by passing to the limit, we see that the action of $\Sigma_{2}^{\prime}$ on ${{\rm Tor}_{1}^{\mathbb{Z}}}(T_{\mathcal{A}},T_{\mathcal{A}})$ is trivial. For the last part note that since $T_{\mathcal{A}}$ is ind-cyclic, $\bigwedge_{\mathbb{Z}}^{3}T_{\mathcal{A}}=0$. Now applying the first part to the inclusion $T_{\mathcal{A}}\hookrightarrow\mathcal{A}$, we obtain the following commutative diagram with exact rows
[…]
Now from this diagram we obtain a natural splitting map
${{\rm Tor}_{1}^{\mathbb{Z}}}(T_{\mathcal{A}},T_{\mathcal{A}})\simeq H_{3}(T_{\mathcal{A}},\mathbb{Z})\rightarrow H_{3}(\mathcal{A},\mathbb{Z}).$
Corollary 1.4.
For any integer $m\in\mathbb{Z}$, let $m:\mathcal{A}\rightarrow\mathcal{A}$ be given by $a\mapsto ma$. Then the map $m_{\ast}:H_{3}(\mathcal{A},\mathbb{Z})\rightarrow H_{3}(\mathcal{A},\mathbb{Z})$ induces multiplication by $m^{3}$ on $\bigwedge_{\mathbb{Z}}^{3}\mathcal{A}$ and multiplication by $m^{2}$ on ${{\rm Tor}_{1}^{\mathbb{Z}}}(\mathcal{A},\mathcal{A})^{\Sigma_{2}^{\prime}}$.
Proof.
This follows from the above proposition. ∎
2. The bloch-wigner exact sequence
Let $A$ be a commutative ring and set $\mathcal{W}_{A}:=\{a\in A:a(1-a)\in{A^{\times}}\}$. The scissors congruence group $\mathcal{P}(A)$ of $A$ is defined as the quotient of the free abelian group generated by symbols $[a]$, $a\in\mathcal{W}_{A}$, by the subgroup generated by the elements
$[a]-[b]+\bigg{[}\frac{b}{a}\bigg{]}-\bigg{[}\frac{1-a^{-1}}{1-b^{-1}}\bigg{]}+\bigg{[}\frac{1-a}{1-b}\bigg{]},$
where $a,b,a/b\in\mathcal{W}_{A}$. Let $S_{\mathbb{Z}}^{2}({A^{\times}}):=({A^{\times}}\otimes_{\mathbb{Z}}{A^{\times}})/\langle a\otimes b+b\otimes a:a,b\in{A^{\times}}\rangle$. The map
$\lambda:\mathcal{P}(A)\rightarrow S_{\mathbb{Z}}^{2}({A^{\times}}),\ \ \ \ [a]\mapsto a\otimes(1-a),$
is well-defined and its kernel is called the Bloch group of $A$ and is denoted by $\mathcal{B}(A)$.
For any local ring $A$ there is a natural map of graded rings $K_{\bullet}^{M}(A)\to K_{\bullet}(A)$, from Milnor $K$-theory to $K$-theory. It is easy to see that
$K_{0}^{M}(A)\simeq K_{0}(A)\simeq\mathbb{Z},\ \ \ \ K_{1}^{M}(A)\simeq K_{1}(A)\simeq{A^{\times}}.$
Moreover, Matsumoto proved that for a field $F$
$K_{2}(F)\simeq K_{2}^{M}(F)\simeq({A^{\times}}\otimes_{\mathbb{Z}}{A^{\times}})/\langle a\otimes(1-a):a\in\mathcal{W}_{F}\rangle$
(see [23, Theorem 4.3.15]).
The indecomposable $K_{3}^{\rm ind}(F)$ of a field $F$ is defined as follows:
$K_{3}^{\rm ind}(F):=K_{3}(F)/K_{3}^{M}(F).$
The Bloch group and the indecomposable part of the third $K$-group are deeply connected.
Theorem 2.1 (Bloch-Wigner exact sequence).
For any field $F$, we have a natural exact sequence
$0\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(F),\mu(F))^{\sim}\rightarrow K_{3}^{\rm ind}(F)\rightarrow\mathcal{B}(F)\rightarrow 0$
Proof.
The case of infinite fields has been proved in [27, Theorem 5.2] and the case of finite fields has been settled in [15, Corollary 7.5]. ∎
For any filed $F$, there are natural maps $H_{2}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z})\rightarrow K_{2}(F)$ and $H_{3}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z})\rightarrow K_{3}^{\rm ind}(F)$ (see [15, §5]).
Proposition 2.2.
If $F$ is a quadratically closed field, then we have natural isomorphisms
$K_{2}(F)\simeq K_{2}^{M}(F)\simeq H_{2}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z}),\ \ \ \ \ K_{3}^{\rm ind}(F)\simeq H_{3}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z}).$
Furthermore, $K_{2}(F)$ is uniquely $2$-divisible.
Proof.
See [24, page 286, 2.12], [17, Proposition 6.4] and [3, Proposition 1.2]. ∎
The following classical Bloch-Wigner exact sequence will be used in the proof of Theorem 3.1.
Corollary 2.3 (Bloch-Wigner, Dupont-Sah).
If $F$ is a quadritically closed field, then we have the exact sequence
${0\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(F),\mu(F))\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z})\rightarrow\mathcal{B}(F)\rightarrow 0.}$
Proof.
For a direct proof see [10, Appedix A]. See also [27, Theorem 5.2]. ∎
Note that the map $\mu(F)\rightarrow\mathit{{\rm SL}}_{2}(F)$ given by $a\mapsto{\left(\!\!\begin{array}[]{cc}\!\!a&\!\!0\\ \!\!0&\!\!a^{-1}\end{array}\!\!\right)}$ induces the natural map
$H_{3}(\mu(F),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z}).$
By Proposition 1.3, $H_{3}(\mu(F),\mathbb{Z})\simeq{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(F),\mu(F))$. Now the left map in the above classical Bloch-Wigner exact sequence is the map
${{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(F),\mu(F))\simeq H_{3}(\mu(F),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z}).$
Proposition 2.4 (Coronado [7]).
Let $F$ be a real closed field. Then we have the Bloch-Wigner exact sequence
${0\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(F),\mu(F))^{\sim}\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z})\rightarrow\mathcal{B}(F)\rightarrow 0.}$
In particular, $H_{3}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z})\simeq K_{3}^{\rm ind}(F)$
Proof.
See [7, Theorem 6.1] and its proof. ∎
3. Low dimensional homology of $\mathit{{\rm SL}}_{2}$ and $\mathit{{\rm PSL}}_{2}$
We apply Theorem 1.1 to the central extension
[…]
where $\mu_{2}(A)=\{a\in{A^{\times}}:a^{2}=1\}$ and the map $i$ is given by $a\mapsto{\left(\!\!\begin{array}[]{cc}\!\!a&\!\!0\\ \!\!0&\!\!a\end{array}\!\!\right)}={\left(\!\!\begin{array}[]{cc}\!\!a&\!\!0\\ \!\!0&\!\!a^{-1}\end{array}\!\!\right)}$. Let
$\widetilde{H}_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z}):=\displaystyle\frac{H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}{\rho_{\ast}\Big{(}\mu_{2}(A)\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\oplus{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu_{2}(A),H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z}))\Big{)}},$
where $\rho:\mu_{2}(A)\times\mathit{{\rm SL}}_{2}(A)\rightarrow\mathit{{\rm SL}}_{2}(A)$ is the product map and is given by $(a,X)\mapsto aX$. We recall that for a field $F$,
$\mu_{2}(F)^{\sim}\simeq\begin{cases}0&\text{if ${\rm char}(F)=2$}\\ \mathbb{Z}/4&\text{if ${\rm char}(F)\neq 2$}.\end{cases}$
Now we are ready to state and prove our first main result.
Theorem 3.1.
Let $A$ be a ring such that there is a ring homomorphism $A\rightarrow F$, $F$ a field, where $\mu_{2}(A)\simeq\mu_{2}(F)$. Then we have the $9$-term exact sequence
$0\rightarrow\mu_{2}(A)^{\sim}\rightarrow\widetilde{H}_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})\rightarrow\mu_{2}(A)\otimes_{\mathbb{Z}}H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$
$\rightarrow H_{2}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\rightarrow H_{2}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})\rightarrow\mu_{2}(A)\rightarrow H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$
$\rightarrow H_{1}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})\rightarrow 0.$
Moreover, if ${\rm char}(F)\neq 2$, then the image of the maps $\mu_{2}(A)^{\sim}\rightarrow\widetilde{H}_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$ and $H_{3}(\langle w\rangle,\mathbb{Z})\rightarrow\widetilde{H}_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$ coincide, where $w:={\left(\!\!\begin{array}[]{cc}\!\!0&\!\!1\\ \!\!-1&\!\!0\end{array}\!\!\right)}$.
Proof.
First we study the homomophism $\tau:\Gamma(\mu_{2}(A))\rightarrow\mu_{2}(A)\otimes_{\mathbb{Z}}H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$ (see (1.1)). Note that by Lemma 1.2, $\Gamma(\mu_{2}(A))\simeq\mu_{2}(A)^{\sim}$. The map $\tau$ is induced by the quadratic map
$\mu_{2}(A)\rightarrow\mu_{2}(A)\otimes_{\mathbb{Z}}H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z}),\ \ \ \ -1\mapsto(-1)\otimes i_{\ast}(-1),$
where $i_{\ast}:\mu_{2}(A)\simeq H_{1}(\mu_{2}(A),\mathbb{Z})\rightarrow H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$ is given by $-1\mapsto{\boldsymbol{c}}(-I_{2})$. Since $w^{2}=-I_{2}$, we have
$(-1)\otimes i_{\ast}(-1)=(-1)\otimes{\boldsymbol{c}}(-I_{2})=(-1)\otimes{\boldsymbol{c}}(w^{2})=(-1)\otimes 2{\boldsymbol{c}}(w)=(-1)^{2}\otimes{\boldsymbol{c}}(w)=0.$
Thus $\tau$ must be trivial. Now by Theorem 1.1 we obtain the exact sequence
$H_{4}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})\rightarrow\Gamma(\mu_{2}(A))\rightarrow\widetilde{H}_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})$
$\rightarrow\mu_{2}(A)\otimes_{\mathbb{Z}}H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\rightarrow\cdots.$
Therefore to complete the proof of the first part of the theorem, we must prove that the map
$H_{4}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})\rightarrow\Gamma(\mu_{2}(A))\simeq\mu_{2}(A)^{\sim}$
is trivial. Since this map is trivial when $\mu_{2}(A)\simeq\mu_{2}(F)=1$, we may assume that $\mu_{2}(A)\simeq\mu_{2}(F)=\{\pm 1\}\simeq\mathbb{Z}/2$.
Let $\overline{F}$ be the algebraic closure of $F$. Then $H_{1}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z})=0$ and $H_{2}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z})$ is uniquely 2-divisible by Proposition 2.2. These imply that ${{\rm Tor}_{1}^{\mathbb{Z}}}(\mu_{2}(\overline{F}),H_{1}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z}))$ and $\mu_{2}(\overline{F})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z})$ are trivial. Hence by Theorem 1.1, we have the exact sequence
(3.1) $\Gamma(\mu_{2}(\overline{F}))\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(\overline{F}),\mathbb{Z})\rightarrow 0.$
Observe that $\Gamma(\mu_{2}(\overline{F}))\simeq\mu_{2}(\overline{F})^{\sim}\simeq\mathbb{Z}/4$ (Lemma 1.2). The surjective map $p:\mathit{{\rm SL}}_{2}(\overline{F})\rightarrow\mathit{{\rm PSL}}_{2}(\overline{F})$ induces the commutative diagram with exact rows
00{0}${{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(\overline{F}),\mu(\overline{F}))}$${H_{3}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z})}$${\mathcal{\mathcal{B}}(\overline{F})}$00{0}00{0}${{{\rm Tor}_{1}^{\mathbb{Z}}}(\widetilde{\mu}(\overline{F}),\widetilde{\mu}(\overline{F}))}$${H_{3}(\mathit{{\rm PSL}}_{2}(\overline{F}),\mathbb{Z})}$${\mathcal{B}(\overline{F})}$${0.}$$\scriptstyle{\alpha}$$\scriptstyle{p_{\ast}}$
The first row is the classical Bloch-Wigner exact sequence discussed in the previous section (Corollary 2.3). The second exact sequence can be proved as the first exact sequence, where $\widetilde{\mu}(\overline{F}):=\mu(\overline{F})/\mu_{2}(\overline{F})$ (see [10, Theorem 4.10 and Appendix A]. Also see [20, Corollary 3.5] below). The left hand side square is obtained from the commutative diagram
${{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(\overline{F}),\mu(\overline{F}))}$${H_{3}(\mu(\overline{F}),\mathbb{Z})}$${H_{3}({\rm T}(\overline{F}),\mathbb{Z})}$${H_{3}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z})}$${{{\rm Tor}_{1}^{\mathbb{Z}}}(\widetilde{\mu}(\overline{F}),\widetilde{\mu}(\overline{F}))}$${H_{3}(\widetilde{\mu}(\overline{F}),\mathbb{Z})}$${H_{3}(\mathit{{\rm PT}}(\overline{F}),\mathbb{Z})}$${H_{3}(\mathit{{\rm PSL}}_{2}(\overline{F}),\mathbb{Z}),}$$\scriptstyle{\alpha}$$\scriptstyle{p_{\ast}}$$\scriptstyle{p_{\ast}}$$\scriptstyle{p_{\ast}}$
where for a ring $R$,
${\rm T}(R):=\bigg{\{}{\left(\!\!\begin{array}[]{cc}\!\!a&\!\!0\\ \!\!0&\!\!a^{-1}\end{array}\!\!\right)}|a\in R^{\times}\bigg{\}},\ \ \ \ \ \mathit{{\rm PT}}(R):={\rm T}(R)/\mu_{2}(R)I_{2}$
and the middle square is induced by the map $\mu(R)\rightarrow{\rm T}(R)$, $a\mapsto{\left(\!\!\begin{array}[]{cc}\!\!a&\!\!0\\ \!\!0&\!\!a^{-1}\end{array}\!\!\right)}$. The extension
$1\rightarrow\mu_{2}(\overline{F})\rightarrow{\rm T}(\overline{F})\rightarrow\mathit{{\rm PT}}(\overline{F})\rightarrow 1$
is isomorphic with the extension
$1\rightarrow\mu_{2}(\overline{F})\rightarrow\overline{F}^{\times}\overset{()^{2}}{\longrightarrow}\overline{F}^{\times}\rightarrow 1$
The map $\phi=()^{2}:\overline{F}^{\times}\rightarrow\overline{F}^{\times}$ induces the map $\phi_{\ast}:H_{3}(\overline{F}^{\times},\mathbb{Z})\rightarrow H_{3}(\overline{F}^{\times},\mathbb{Z})$. By Proposition 1.3 we have the split exact sequence
$\begin{array}[]{c}0\rightarrow\bigwedge_{\mathbb{Z}}^{3}\overline{F}^{\times}\rightarrow H_{3}(\overline{F}^{\times},\mathbb{Z})\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(\overline{F}),\mu(\overline{F}))\rightarrow 0.\end{array}$
By Corollary 1.4 this leads to multiplication by $2^{2}$ on ${{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(\overline{F}),\mu(\overline{F}))$. Since
${{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(\overline{F}),\mu(\overline{F}))\simeq\mu(\overline{F}),$
the above fact implies that $\ker(\alpha)\simeq\mathbb{Z}/4$. On the other hand by the Snake lemma $\ker(\alpha)\simeq\ker(p_{\ast})$. Thus the left map of the exact sequence (3.1) must be injective and so we have the exact sequence
$0\rightarrow\Gamma(\mu_{2}(\overline{F}))\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(\overline{F}),\mathbb{Z})\rightarrow 0$
Observe that
$\Gamma(\mu_{2}(\overline{F}))\simeq\mu_{2}(\overline{F})^{\sim}\hookrightarrow\mu(\overline{F})^{\sim}\simeq\mu(\overline{F})\simeq{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(\overline{F}),\mu(\overline{F})).$
The composite $A\rightarrow F\hookrightarrow\overline{F}$ induces the morphism of extensions
[…]
Now by Theorem 1.1 we obtain the commutative diagram with exact rows
${H_{4}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})}$${\Gamma(\mu_{2}(A))}$${\widetilde{H}_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}$${H_{3}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})}$00{0}${\Gamma(\mu_{2}(\overline{F}))}$${H_{3}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z})}$${H_{3}(\mathit{{\rm PSL}}_{2}(\overline{F}),\mathbb{Z})}$${0,}$
where the exactness of the second row has just been shown. Since $\mu_{2}(A)\simeq\mu_{2}(F)=\mu_{2}(\overline{F})$, we have
$\Gamma(\mu_{2}(A))\simeq\Gamma(\mu_{2}(\overline{F})).$
Now from the above diagram it follows that the map $\Gamma(\mu_{2}(A))\rightarrow\widetilde{H}_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$ is injective. This complete the proof of the first claim.
For the second claim, first we show that the composite
$H_{3}(\langle w\rangle,\mathbb{Z})\rightarrow\widetilde{H}_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})$
is trivial. Let denote this map by $\varphi_{\ast}$. The group $H_{3}(\langle w\rangle,\mathbb{Z})$ is cyclic of order 4 and is generated by
$\chi:=\overline{[w|w|w]+[w|\!-\!1|w]+[w|\!-\!w|w]}\in H_{3}(B_{\bullet}(\langle w\rangle)_{\langle w\rangle})=H_{3}(\langle w\rangle,\mathbb{Z}),$
(see [29, Proposition 8.1] or [22, Proposition 3.25]), where $B_{\bullet}(\langle w\rangle)\rightarrow\mathbb{Z}$ is the bar resolution of $\mathbb{Z}$ over $\langle w\rangle$. We have
$\varphi_{\ast}(\chi)=\overline{[w|w|w]+[w|\!-\!w|w]}=\overline{2[w|w|w]}.$
Since
$d_{4}([w|w|w|w])={2[w|w|w]}$
in $B_{\bullet}(\mathit{{\rm PSL}}_{2}(A))_{\mathit{{\rm PSL}}_{2}(A)}$, we have $\varphi_{\ast}(\chi)=0$. This shows that the above composite is trivial. Now by the commutative diagram
${H_{3}(\langle w\rangle,\mathbb{Z})}$${\widetilde{H}_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}$${H_{3}(\langle w\rangle,\mathbb{Z})}$${H_{3}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z}),}$
it is sufficient to prove that the map $H_{3}(\langle w\rangle,\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z})$ is injective. Let $\mathfrak{i}^{2}=-1$. Then $w$ is conjugate to ${\left(\!\!\begin{array}[]{cc}\!\!\mathfrak{i}&\!\!0\\ \!\!0&\!\!\mathfrak{i}^{-1}\end{array}\!\!\right)}$. It is known by the classical Bloch-Wigner exact sequence (Corollary 2.3) that the map
$H_{3}(\mu(\overline{F}),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z})$
which is induced by $\mu(\overline{F})\rightarrow\mathit{{\rm SL}}_{2}(\overline{F})$, $a\mapsto{\left(\!\!\begin{array}[]{cc}\!\!a&\!\!0\\ \!\!0&\!\!a^{-1}\end{array}\!\!\right)}$, is injective. Thus the map
$H_{3}(\langle\mathfrak{i}\rangle,\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z})$
is injective. This proves our claim. ∎
Remark 3.2.
The above theorem is also valid if we replace $\mathit{{\rm SL}}_{2}(A)$ with the elementary subgroup ${\rm E}_{2}(A)$. The proof stays the same. Just one should note that over a field $F$, ${\rm E}_{2}(F)=\mathit{{\rm SL}}_{2}(F)$.
Corollary 3.3.
Let $A$ be a ring such that there is a ring homomorphism $A\rightarrow F$, $F$ a field of ${\rm char}(F)\neq 2$, where $\mu_{2}(A)\simeq\mu_{2}(F)$. Then the map $H_{3}(\langle w\rangle,\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$ is injective.
Proof.
By Theorem 3.1, the map $H_{3}(\langle w\rangle,\mathbb{Z})\rightarrow\widetilde{H}_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$ is injective. But this map factors through $H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$. Thus we have our claim. ∎
4. The low dimensional homology of $\mathit{{\rm SL}}_{2}(\mathbb{Z}[\frac{1}{m}])$ and $\mathit{{\rm PSL}}_{2}(\mathbb{Z}[\frac{1}{m}])$
It is known that $\mathit{{\rm SL}}_{2}(\mathbb{Z})\simeq\mathbb{Z}/4\ast_{\mathbb{Z}/2}\mathbb{Z}/6$ and $\mathit{{\rm PSL}}_{2}(\mathbb{Z})\simeq\mathbb{Z}/2\ast\mathbb{Z}/3$. By the Mayer-Vietoris exact sequence associated to a amalgamated product [6, Corollary 7.7, Chap. II] one obtains
$H_{1}(\mathit{{\rm SL}}_{2}(\mathbb{Z}),\mathbb{Z})\simeq\mathbb{Z}/12,\ \ \ \ \ H_{2}(\mathit{{\rm SL}}_{2}(\mathbb{Z}),\mathbb{Z})=0,$
$H_{3}(\mathit{{\rm SL}}_{2}(\mathbb{Z}),\mathbb{Z})\simeq\mathbb{Z}/12,\ \ \ H_{3}(\mathit{{\rm PSL}}_{2}(\mathbb{Z}),\mathbb{Z})\simeq\mathbb{Z}/6.$
Using these and Theorem 3.1 we obtain the exact sequence
(4.1) $0\rightarrow\mathbb{Z}/4\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(\mathbb{Z}),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(\mathbb{Z}),\mathbb{Z})\rightarrow\mathbb{Z}/2\rightarrow 0,$
which will be used in the proof of the following result.
Proposition 4.1.
Let $m$ be a square free integer and let $A_{m}:=\mathbb{Z}[\frac{1}{m}]$.
(i) If $m$ is even, then $H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})\simeq H_{1}(\mathit{{\rm PSL}}_{2}(A_{m}),\mathbb{Z})$ and the sequences
$0\rightarrow H_{2}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})\rightarrow H_{2}(\mathit{{\rm PSL}}_{2}(A_{m}),\mathbb{Z})\rightarrow\mu_{2}(A_{m})\rightarrow 0,$
$\begin{array}[]{c}0\rightarrow\mathbb{Z}/4\rightarrow{\displaystyle\frac{H_{3}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})}{\mu_{2}(A_{m})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})}}\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(A_{m}),\mathbb{Z})\rightarrow 0\end{array}$
are exact.
(ii) If $m$ is odd, then $H_{2}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})\simeq H_{2}(\mathit{{\rm PSL}}_{2}(A_{m}),\mathbb{Z})$ and the sequences
$0\rightarrow\mu_{2}(A_{m})\rightarrow H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})\rightarrow H_{1}(\mathit{{\rm PSL}}_{2}(A_{m}),\mathbb{Z})\rightarrow 0,$
$\begin{array}[]{c}0\rightarrow\mathbb{Z}/4\rightarrow{\displaystyle\frac{H_{3}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})}{\mu_{2}(A_{m})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})}}\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(A_{m}),\mathbb{Z})\rightarrow\mathbb{Z}/2\rightarrow 0\end{array}$
are exact.
Proof.
By [19, Proposition 4.3] or [21, Theorem 1.2] we have
$H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})\simeq\begin{cases}0&\text{if $2\mid m$, $3\mid m$}\\ \mathbb{Z}/3&\text{if $2\mid m$, $3\nmid m$}\\ \mathbb{Z}/4&\text{if $2\nmid m$, $3\mid m$.}\\ \mathbb{Z}/12&\text{if $2\nmid m$, $3\nmid m$}\end{cases}$
It follows from this that the inclusion $\mathbb{Z}\subseteq A_{m}$ induces the surjective splitting map
$\mathbb{Z}/12\simeq H_{1}(\mathit{{\rm SL}}_{2}(\mathbb{Z}),\mathbb{Z})\twoheadrightarrow H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z}).$
Now from the commutative diagram
${{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu_{2}(\mathbb{Z}),H_{1}(\mathit{{\rm SL}}_{2}(\mathbb{Z}),\mathbb{Z}))}$${H_{3}(\mathit{{\rm SL}}_{2}(\mathbb{Z}),\mathbb{Z})}$${{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu_{2}(A_{m}),H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z}))}$${\displaystyle\frac{H_{3}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})}{\mu_{2}(A_{m})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})}}$$\scriptstyle{\rho_{\ast}=0}$$\scriptstyle{\rho_{\ast}}$
it follows that the bottom map is trivial. In the above diagram, first note that since $H_{2}(\mathit{{\rm SL}}_{2}(\mathbb{Z}),\mathbb{Z})=0$, we have $H_{3}(\mathit{{\rm SL}}_{2}(\mathbb{Z}),\mathbb{Z})=\displaystyle\frac{H_{3}(\mathit{{\rm SL}}_{2}(\mathbb{Z}),\mathbb{Z})}{\mu_{2}(\mathbb{Z})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(\mathbb{Z}),\mathbb{Z})}$ and thus the above diagram is commutative. Second, by the exact sequence (4.1), the upper map in the above diagram should be trivial.
(i) Let $m$ be even. Then from the above calculation of $H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})$, we see that
$\mu_{2}(A_{m})\otimes H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})=0$
and the natural map $\mu_{2}(A_{m})\rightarrow H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})$ is trivial. Thus the claim follows from Theorem 3.1.
(ii) Now assume that $m$ is odd. Then $\mu_{2}(A_{m})\otimes H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})\simeq\mathbb{Z}/2$. Moreover, from the commutative diagram
${H_{3}(\mathit{{\rm PSL}}_{2}(\mathbb{Z}),\mathbb{Z})}$${\mu_{2}(\mathbb{Z})\otimes_{\mathbb{Z}}H_{1}(\mathit{{\rm SL}}_{2}(\mathbb{Z}),\mathbb{Z})}$${H_{3}(\mathit{{\rm PSL}}_{2}(A_{m}),\mathbb{Z})}$${\mu_{2}(A_{m})\otimes_{\mathbb{Z}}H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})}$
it follows that the bottom map is surjective (note that the upper map is surjective in view of Theorem 3.1, because $H_{2}(\mathit{{\rm SL}}_{2}(\mathbb{Z}),\mathbb{Z})=0$). Since
$H_{1}(\mathit{{\rm PSL}}_{2}(A_{m}),\mathbb{Z})\simeq(\mathbb{Z}/2)^{\alpha}\oplus(\mathbb{Z}/3)^{\beta}$
for some $\alpha$ and $\beta$ (see [28, Corollary 4.4]), from the group structure of $H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})$ one sees that the natural map
$\mu_{2}(A_{m})\rightarrow H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})$
must be injective. Hence the claim follows from Theorem 3.1. ∎
Remark 4.2.
A central extension $1\rightarrow\mathcal{A}\rightarrow\mathcal{G}\rightarrow\mathcal{Q}\rightarrow 1$ is called a stem-extension if the natural map
$\mathcal{A}=H_{1}(\mathcal{A},\mathbb{Z})\rightarrow H_{1}(\mathcal{G},\mathbb{Z})=\mathcal{G}^{\rm ab}$
is trivial [13, §4] and it is called a weak stem-extension if the natural map
$\mathcal{A}\otimes_{\mathbb{Z}}\mathcal{A}\rightarrow\mathcal{A}\otimes_{\mathbb{Z}}H_{1}(\mathcal{G},\mathbb{Z})$
is trivial [13, page 103]. The proof of Theorem 3.1 shows that if $A$ is a ring such that $\mu_{2}(A)=\{\pm 1\}$, then
$1\rightarrow\mu_{2}(A)\rightarrow\mathit{{\rm SL}}_{2}(A)\rightarrow\mathit{{\rm PSL}}_{2}(A)\rightarrow 1$
always is a weak stem-extension. But by Proposition 4.1 this extension is not a stem-extension for $A=\mathbb{Z}[\frac{1}{m}]$, where $m$ is a square free odd integer. This also is true for $A=\mathbb{Z}$ by the paragraph above Proposition 4.1 and $A=\mathbb{Z}_{(2)}$ by Example 5.3.
We learned the following result from Kevin Hutchinson. It seems that this lemma is widely known, but we could not find any suitable reference for it.
Lemma 4.3.
Let $\mathcal{G}$ be a group such that $H_{n}(\mathcal{G},\mathbb{Z})$, $H_{n+1}(\mathcal{G},\mathbb{Z})$ and $H^{n+1}(\mathcal{G},\mathbb{Z})$ are finitely generated. Then
${\rm rank}\ H_{n+1}(\mathcal{G},\mathbb{Z})={\rm rank}\ H^{n+1}(\mathcal{G},\mathbb{Z}),\ \ \ and\ \ \ H_{n}(\mathcal{G},\mathbb{Z})_{\rm tor}\simeq H^{n+1}(\mathcal{G},\mathbb{Z})_{\rm tor}.$
Proof.
This claim easily follows from the Universal Coefficients Theorem
$0\rightarrow{{\rm Ext}_{\mathbb{Z}}^{1}}(H_{n}(\mathcal{G},\mathbb{Z}),\mathbb{Z})\rightarrow H^{n+1}(\mathcal{G},\mathbb{Z})\rightarrow{\rm Hom}(H_{n+1}(\mathcal{G},\mathbb{Z}),\mathbb{Z})\rightarrow 0$
and the facts that for any abelian group $\mathcal{A}$, ${{\rm Ext}_{\mathbb{Z}}^{1}}(\mathbb{Z},\mathcal{A})=0$ and ${{\rm Ext}_{\mathbb{Z}}^{1}}(\mathbb{Z}/n,\mathcal{A})\simeq\mathcal{A}/n\mathcal{A}$. ∎
Example 4.4.
Let $A_{m}:=\mathbb{Z}[\frac{1}{m}]$, where $m=2$ or $3$.
(i) The product map $\mu_{2}(A_{2})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A_{2}),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(A_{2}),\mathbb{Z})$ is trivial: By [20, Corolllary 3.7, Remark 3.8] we have the isomorphism
$H_{3}(\mathit{{\rm PSL}}_{2}(A_{2}),\mathbb{Z})\simeq\mathcal{RB}(A_{2}),$
where $\mathcal{RB}(A_{2})$ is the refined Bloch group of $A_{2}$ (see [8, §4] for the definition of $\mathcal{RB}(A_{2})$). By a result of Coronado and Hutchinson, we have the refined Bloch-Wigner exact sequence
$0\rightarrow\mathbb{Z}/4\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(A_{2}),\mathbb{Z})\rightarrow\mathcal{RB}(A_{2})\rightarrow 0$
(see [8, Proposition 8.31]). From these two results and Proposition 4.1, it follows that the desired product map is trivial.
(ii) Now we show that the product map $\mu_{2}(A_{3})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A_{3}),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(A_{3}),\mathbb{Z})$ is not trivial. We know that
$H_{3}(\mathit{{\rm SL}}_{2}(A_{3}),\mathbb{Z})\simeq\mathbb{Z}/12\oplus\mathbb{Z}/4.$
This follows from the cohomology calculation of $\mathit{{\rm SL}}_{2}(A_{3})$ in [1, page 8] and Lemma 4.3. Furthermore, by [28, Proposition 4.2], $H^{k}(\mathit{{\rm PSL}}_{2}(A_{3}),\mathbb{Z})$ is generated by elements of orders $2$ and $3$ for $k\geq 3$. Again by Lemma 4.3 it follows that $H_{3}(\mathit{{\rm PSL}}_{2}(A_{3}),\mathbb{Z})$ is generated by elements of orders $2$ and $3$. Now let the above product map be trivial. Then by Proposition 4.1 the sequence
$\begin{array}[]{c}0\rightarrow\mathbb{Z}/4\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(A_{3}),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(A_{3}),\mathbb{Z})\rightarrow\mathbb{Z}/2\rightarrow 0\end{array}$
must be exact. This and the isomorphism $H_{3}(\mathit{{\rm SL}}_{2}(A_{3}),\mathbb{Z})\simeq\mathbb{Z}/12\oplus\mathbb{Z}/4$ imply that $H_{3}(\mathit{{\rm PSL}}_{2}(A_{3}),\mathbb{Z})$ has an element of order $4$, which is a contradiction. Observe that $H_{2}(\mathit{{\rm SL}}_{2}(A_{3}),\mathbb{Z})\simeq\mathbb{Z}$ and thus $\mu_{2}(A_{3})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A_{3}),\mathbb{Z})\simeq\mathbb{Z}/2$.
5. The low dimensional homology groups of $\mathit{{\rm SL}}_{2}$ and $\mathit{{\rm PSL}}_{2}$ of local rings
Let $A$ be a local ring with maximal ideal $\mathfrak{m}_{A}$. It is known that
$H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\simeq\begin{cases}A/\mathfrak{m}_{A}^{2}&\text{if $|A/\mathfrak{m}_{A}|=2$}\\ A/\mathfrak{m}_{A}&\text{if $|A/\mathfrak{m}_{A}|=3$.}\\ 0&\text{if $|A/\mathfrak{m}_{A}|\geq 4$}\\ \end{cases}$
(see [19, Proposition 4.1]). Now let $|A/\mathfrak{m}_{A}|\neq 2$. Since $\mu_{2}(A)\otimes_{\mathbb{Z}}H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})=0$, from Theorem 1.1 we obtain the exact sequence
$0\rightarrow H_{2}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\rightarrow H_{2}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})\rightarrow\mu_{2}(A)\rightarrow 0$
Since
$\mu_{2}(A)\otimes_{\mathbb{Z}}H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})=0,\ \ \ \ {{\rm Tor}_{1}^{\mathbb{Z}}}(\mu_{2}(A),H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z}))=0,$
by Theorem 3.1 we obtain the following result.
Proposition 5.1.
Let $A$ be a local ring such that there is a ring homomorphism $A\rightarrow F$, $F$ a field, where $\mu_{2}(A)\simeq\mu_{2}(F)$. If $|A/\mathfrak{m}_{A}|\neq 2$, then the sequence
$0\rightarrow\mu_{2}(A)^{\sim}\rightarrow\frac{H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}{\mu_{2}(A)\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})\rightarrow 0$
is exact.
Remark 5.2.
The result of the above proposition is valid for any domain $A$ such that $H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$ is torsion of odd exponent. For example let $\mathcal{O}_{-3}$ be the ring of algebraic integers of the imaginary quadratic field $\mathbb{Q}(\sqrt{-3})$. The homology groups of $\mathit{{\rm SL}}_{2}(\mathcal{O}_{-3})$ and $\mathit{{\rm PSL}}_{2}(\mathcal{O}_{-3})$ have been calculated in [2, Theorem, page 376] and [25, Theorem 5.7]. We list these homology groups for $1\leq n\leq 3$:
$H_{n}(\mathit{{\rm SL}}_{2}(\mathcal{O}_{-3}),\mathbb{Z})\simeq\begin{cases}\mathbb{Z}/3&\text{if $n=1$}\\ \mathbb{Z}/4&\text{if $n=2,$}\\ \mathbb{Z}/24\times\mathbb{Z}/6&\text{if $n=3$}\end{cases}$
$H_{n}(\mathit{{\rm PSL}}_{2}(\mathcal{O}_{-3}),\mathbb{Z})\simeq\begin{cases}\mathbb{Z}/3&\text{if $n=1$}\\ \mathbb{Z}/4\times\mathbb{Z}/2&\text{if $n=2.$}\\ \mathbb{Z}/6\times\mathbb{Z}/3&\text{if $n=3$}\end{cases}$
From Theorem 3.1 we have the exact sequence
$\begin{array}[]{c}0\rightarrow\mathbb{Z}/4\rightarrow{\displaystyle\frac{H_{3}(\mathit{{\rm SL}}_{2}(\mathcal{O}_{-3}),\mathbb{Z})}{\mu_{2}(\mathcal{O}_{-3})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(\mathcal{O}_{-3}),\mathbb{Z})}}\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(\mathcal{O}_{-3}),\mathbb{Z})\rightarrow 0.\end{array}$
From the sizes of the third homology groups of $\mathit{{\rm SL}}_{2}(\mathcal{O}_{-3})$ and $\mathit{{\rm PSL}}_{2}(\mathcal{O}_{-3})$ it follows that the product map
$\mu_{2}(\mathcal{O}_{-3})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(\mathcal{O}_{-3}),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(\mathcal{O}_{-3}),\mathbb{Z})$
is non-trivial.
Example 5.3.
Let $A$ be a local domain such that $A/\mathfrak{m}_{A}\simeq\mathbb{F}_{2}$. Then $H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\simeq A/\mathfrak{m}_{A}^{2}$. By a careful analysis of the composition
$\mu_{2}(A)\rightarrow H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\overset{\simeq}{\longrightarrow}A/\mathfrak{m}_{A}^{2},$
one sees that $-1\in\mu_{2}(A)$ maps to $-\overline{6}\in A/\mathfrak{m}_{A}^{2}$. Since $2\in\mathfrak{m}_{A}$, we have $-\overline{6}=\overline{2}$ in $A/\mathfrak{m}_{A}^{2}$. Thus if $2\in\mathfrak{m}_{A}^{2}$, then $H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\simeq H_{1}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})$. Otherwise we have the exact sequence
$1\rightarrow\mu_{2}(A)\rightarrow H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\rightarrow H_{1}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})\rightarrow 0$
Observe that
$\mu_{2}(A)\otimes_{\mathbb{Z}}H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\simeq\mu_{2}(A)\otimes_{\mathbb{Z}}A/\mathfrak{m}_{A}^{2}\simeq A/(2A+\mathfrak{m}_{A}^{2}).$
If $A=\mathbb{Z}_{(2)}=\{a/b\in\mathbb{Q}:a,b\in\mathbb{Z},2\nmid b\}$ (or $A=\mathbb{Z}_{2}$, the ring of $2$-adic integers), then $2\notin\mathfrak{m}_{A}^{2}$ (by the Nakayama lemma). Thus the sequence
$1\rightarrow\mu_{2}(\mathbb{Z}_{(2)})\rightarrow H_{1}(\mathit{{\rm SL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})\rightarrow H_{1}(\mathit{{\rm PSL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})\rightarrow 0$
is exact. Note that $H_{1}(\mathit{{\rm SL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})\simeq\mathbb{Z}/4$ [19, Example 4.2]. With an argument similar to the proof of Proposition 4.1(ii), one can show that
${{\rm Tor}_{1}^{\mathbb{Z}}}(\mu_{2}(\mathbb{Z}_{(2)}),H_{1}(\mathit{{\rm SL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z}))\rightarrow\displaystyle\frac{H_{3}(\mathit{{\rm SL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})}{\mu_{2}(\mathbb{Z}_{(2)})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})}$
is trivial and
$H_{3}(\mathit{{\rm PSL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})\rightarrow\mu_{2}(\mathbb{Z}_{(2)})\otimes_{\mathbb{Z}}H_{1}(\mathit{{\rm SL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})\simeq\mathbb{Z}/2$
is surjective. Thus
$H_{2}(\mathit{{\rm SL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})\simeq H_{2}(\mathit{{\rm PSL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})$
and the sequence
$\begin{array}[]{c}0\rightarrow\mathbb{Z}/4\rightarrow{\displaystyle\frac{H_{3}(\mathit{{\rm SL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})}{\mu_{2}(\mathbb{Z}_{(2)})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})}}\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})\rightarrow\mathbb{Z}/2\rightarrow 0\end{array}$
is exact. We have similar results for $\mathbb{Z}_{2}$. Hence Proposition 5.1 is not necessarily true when $|A/\mathfrak{m}_{A}|=2$.
Proposition 5.4.
Let $F$ be either a finite field, a quadratically closed field, or a real closed field. Then the sequence
$0\rightarrow\mu_{2}(F)^{\sim}\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(F),\mathbb{Z})\rightarrow 0$
is exact. In particular, the product map
$\mu_{2}(F)\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z})$
is trivial.
Proof.
(i) Let $F=\mathbb{F}_{q}$ be finite. For fields of characteristic $2$ the above exact sequence is trivial. Thus we may assume that $q$ is odd. We have seen that
$H_{1}(\mathit{{\rm SL}}_{2}(\mathbb{F}_{q}),\mathbb{Z})\simeq\begin{cases}0&\text{if $q\neq 3$}\\ \mathbb{Z}/3&\text{if $q=3$.}\end{cases}$
Moreover, it is well-known that
$H_{2}(\mathit{{\rm SL}}_{2}(\mathbb{F}_{q}),\mathbb{Z})\simeq\begin{cases}0&\text{if $q\neq 9$}\\ \mathbb{Z}/3&\text{if $q=9$}\end{cases}$
(see [24, page 280]). These facts together with Theorem 3.1 give us the desired exact sequence.
(ii) Let $F$ be quadratically closed. By Proposition 2.2, $H_{2}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z})$ is uniquely 2-divisible. Thus by Proposition 5.1 we obtain the desired exact sequence.
(iii) Let $F$ be a real closed field. By Proposition 2.4 we have the Bloch-Wigner exact sequence
$0\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(F),\mu(F))^{\sim}\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z})\rightarrow\mathcal{B}(F)\rightarrow 0$
From [20, Corollary 3.5] and [7, Corollary 4.7] we obtain the isomorphism
$H_{3}(\mathit{{\rm PSL}}_{2}(F),\mathbb{Z})\simeq\mathcal{B}(F).$
Our claim follows from these two facts. Thus we presented a proof of this in above proposition. ∎
Remark 5.5.
It is mentioned in [11, page 230] that the sequence
$0\rightarrow\mathbb{Z}/4\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(\mathbb{R}),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(\mathbb{R}),\mathbb{Z})\rightarrow 0$
is exact and for a proof the authors referred to [10, App. A]. We couldn’t really verify this. So we gave a proof of this in above proposition.